# Supervised Knowledge Makes Large Language Models Better In-context Learners

### 26 Dec 2023 v1

@@ Abstract
Large Language Models (LLMs) exhibit emerging in-context learning abilities through prompt engineering. The recent progress in large-scale generative models has further expanded their use in real-world language applications. However, the critical challenge of improving the generalizability and factuality of LLMs in natural language understanding and question answering remains under-explored. While previous in-context learning research has focused on enhancing models to adhere to users’ specific instructions and quality expectations, and to avoid undesired outputs, little to no work has explored the use of task-Specific fine-tuned Language Models (SLMs) to improve LLMs’ in-context learning during the inference stage. Our primary contribution is the establishment of a simple yet effective framework that enhances the reliability of LLMs as it: 1) generalizes out-of-distribution data, 2) elucidates how LLMs benefit from discriminative models, and 3) minimizes hallucinations in generative tasks. Using our proposed plug-in method, enhanced versions of Llama 2 and ChatGPT surpass their original versions regarding generalizability and factuality. We offer a comprehensive suite of resources, including 16 curated datasets, prompts, model checkpoints, and LLM outputs across 9 distinct tasks. Our empirical analysis sheds light on the advantages of incorporating discriminative models into LLMs and highlights the potential of our methodology in fostering more reliable LLMs.
@@ 3.1 Setup
Source models. As reported in GLUE-X (Yang et al., 2022), ELECTRA-large (Clark et al., 2020) achieves the best performance for both ID and OOD tasks over 21 small-scale pre-trained language models (maximum 774M parameters). Hence, we select ELECTRA-large as the SLM for NLU experiments, and RoBERTa-large Liu et al. (2019) for QA. For evaluating the performance of SLM-enhanced LLMs, we select ChatGPT (OpenAI, 2023a) and Llama2-7B-chat (Touvron et al., 2023) as backbones, which are pre-trained on CommonCrawl, WebText, English Wiki, and others.
@@ 3.2 NLU Results
We also present the results of SuperContext with the prompt of the interpreter, which requires LLM to recall influential in-context examples and output rationales when making the predictions, indicating as SuperContext (+interpreter). To better understand the benefits of including the model confidence in the prompt, we present the results of SuperContext (w/o confidence). By comparing the results of SuperContext w/ and w/o confidence, we observe that including model confidence can bring significant improvements in the average performance for both ChatGPT and Llama2. Meanwhile, we find that for QNLI and QQP, Llama2 without the SLM’s confidence achieves the best performance among several methods. Our results also indicate that the interpreter can not bring significant benefits when compared to SuperContext in most of the tasks, except a slight improvement can be achieved on QNLI. It can be because the explain-then-predict prompt (Wang et al., 2022) may not be suitable for incorporating with SuperContext, leading to information overload.
@@ 3.3 QA Results
Results of ChatGPT. The results are presented in Table 3. We find that although the original ChatGPT can achieve the highest accuracy for deterministic questions (79.35%), the exact match (EM) and accuracy for open questions can be significantly improved by SuperContext. In particular, the accuracy for open questions can be improved from 32.35% (ChatGPT) to 54.65% (SuperContext), indicating the huge benefits. Besides, we find that even with the careful design of in-context learning prompts and filter methods, SuperContext still outperforms two in-context learning variants in terms of all metrics, indicating that pure in-context learning without fine-tuning LLMs brings no benefit to the QA task. Furthermore, SuperContext even outperforms the fine-tuned method in a multi-turn setting on all metrics. We believe that such a huge performance benefit (54.65% vs. 13.22%) compared to the traditional in-context method proves that results achieved by discriminative models are effective enough to reduce the hallucination in the QA task.
@@ 4.2 Interpretation Analysis
Learning from In-context Demonstrations. We explore how language models use long contexts. Figure 3 shows the influence of demonstrations during the inference stage, where the y-axis indicates how many times ChatGPT and InstructGPT take the $i_{th}$ in-context example as the emphasized one towards the prediction. The x-axis is sorted by the order of occurrence of in-context examples over 8 natural language understanding tasks. As shown in the figure, both ChatGPT and InstructGPT show a significant occurrence times difference among in-context examples. In particular, ChatGPT with 16-shot examples shows a trend of decreasing attention with the order of appearance. For example, the second in-context example has been paid attention to over 35,000 times while the last example only receives around 5,000 times attention. In terms of InstructGPT, we observe distinctive U-shaped occurrence times, which can be visualized in Figure 3(b). We find that the model tends to pay attention to the beginning or the end of the input context (in-context examples), and the attention significantly degrades in the middle of long contexts. This observation is consistent with the findings of Liu et al. (2023) on the use of long contexts when performing downstream tasks, which suggests that model performance significantly degrades when models must access relevant information in the middle of long contexts and provide a new perspective for future long-context models.
@@
Output Rationale. We also collect the sentence-level rationale generated by LLMs when making predictions, and count for the work frequency for each task of GLUE-X based on ChatGPT, aiming to provide the internal causes of OOD generalizability. According to the human evaluation – finished by 3 people with a CS Ph.D. degree from Europe and the U.S. – of the top 20 influential words/phrases in terms of word frequency, we find that SuperContext achieves the highest average rationale overlap between humans and models after removing the stop word by the NLTK package, 0.575, while the 16-shot ChatGPT only achieves 0.45 overlap score. The IAA between the annotators is 0.92, indicating a high confidence level. However, the rationale is generated by LLMs and thus may contain hallucinations, which should be treated with caution and just for reference.
@@ 4.3 The Effect of SLM Confidence
Since we rely on the complementarity between SLMs and PLMs, SLMs must convey its certainly in task knowledge and uncertainly in domain knowledge to PLMs. The confidence score in the design serves a crucial role in such communication channels. We show the correlation between the confidence of SLMs and the prediction performance of LLMs. As shown in Figure 4, both ChatGPT and Llama2-7B-chat demonstrate a positive correlation between SLMs’ confidence and LLM’ performance, representing a high consistency between those models. The x-axis represents the confidence interval covering from 0.4-1.0, for example, 0.5 indicates the instances with the prediction confidence between 0.4-0.5. It is noteworthy that the confidence is computed by the zero-shot test based on SLMs trained on unseen domains, which indicates that high confidence requires a decent generalization ability of small models. We speculate that SuperContext shows superior performance than both SLMs and LLMs since it leverages the benefits of high consistency in discriminative models and the complementarity property of recent generative models. Besides, such a positive calibration law underscores the importance of involving both prediction and confidence in the prompt design of SuperContext. The data statistic of data quantity shows that most instances included in GLUE-X receive the highest confidence interval from 0.9 to 1.0, and this part of the data can be predicted with significantly higher accuracy than others. By comparing the experimental results of GPT-3.5 and Llama2-7B-chat, we find that there is no significant performance gap between these models for those instances with the confidence of SLMs less than 0.6. When the confidence is more than 0.6, the average performance of GPT-3.5 is substantially better than Llama2-7B-chat. It can be because LLMs with larger parameters are more sensitive to numbers (confidence score) and possibly more faithful to the SLMs’ reference.
@@ 5 Related Work
There is a line of work for improving the in-context learning performance by either constructing demonstrations (Arora et al., 2022; Si et al., 2022; Lyu et al., 2022; Gu et al., 2023; Ye et al., 2023; Dhuliawala et al., 2023) or framing an exploration of example selection methods (Wu et al., 2023; Wang et al., 2023b; Agrawal et al., 2022; Sun et al., 2023a) and even order (Lu et al., 2021; Zhao et al., 2021; Liu et al., 2021; 2023). The contrastive decoding method (Li et al., 2023b) considers the assistance smaller language model but requires external computation. Differently, SuperContext demonstrates its superior performance on OOD test data in a cost-effective manner.
@@
Knowledge in Context. Using external knowledge as auxiliary information to assist LLMs in providing truthful and timely responses represents an emerging solution (Mialon et al., 2023) in recent. Traditional retrieve-based methods (Rubin et al., 2021; Ni et al., 2021; King & Flanigan, 2023) require a knowledge retriever as the prior step for guiding the generation of responses. Besides, the external knowledge source could extend beyond local documents to encompass the entire Internet (Ni et al., 2021; Gao et al., 2023). In addition, LLMs can leverage special plug-ins to improve their capabilities, such as Toolformer (Schick et al., 2023) and LangChain (Chase, 2022) for calling external APIs, and HuggingGPT (Shen et al., 2023) to complete the planning task.
Previous work either relies on web information and search engines for gaining external knowledge (Yu et al., 2023) or accomplishes planning tasks outside the NLP scope. (Xu et al., 2023a) evaluates the efficacy of language model plug-ins under an in-domain setting and lacks an interpretation part to explain the reasons. SuperContext shares a conceptual similarity with HuggingGPT in leveraging Language Model architectures. However, the key distinction lies in our approach’s application and analysis under out-of-distribution (OOD) conditions, a less explored area in the existing literature.
@@ Appendix C Additional Results of Calibration Laws
We supplement the results of the calibration laws with two additional model groups, namely ELECTRA-large and InstrutGPT, in Figure 5. Consistent with ChatGPT and Llama2-7B-chat, both ELECTRA-large and InstruGPT exhibit a positive correlation between confidence and performance. Distinctively, the curve for InstrutGPT demonstrates more pronounced fluctuations, especially when the confidence is relatively low.

### 11 Apr 2024 v2

@@ Abstract
Large Language Models (LLMs) exhibit emerging in-context learning abilities through prompt engineering. The recent progress in large-scale generative models has further expanded their use in real-world language applications. However, the critical challenge of improving the generalizability and factuality of LLMs in natural language understanding and question answering remains under-explored. While previous in-context learning research has focused on enhancing models to adhere to users’ specific instructions and quality expectations, and to avoid undesired outputs, little to no work has explored the use of task-Specific fine-tuned Language Models (SLMs) to improve LLMs’ in-context learning during the inference stage. Our primary contribution is the establishment of a simple yet effective framework that enhances the reliability of LLMs as it: 1) generalizes out-of-distribution data, 2) elucidates how LLMs benefit from discriminative models, and 3) minimizes hallucinations in generative tasks. Using our proposed plug-in method, enhanced versions of Llama 2 and ChatGPT surpass their original versions regarding generalizability and factuality. We offer a comprehensive suite of resources, including 16 curated datasets, prompts, model checkpoints, and LLM outputs across 9 distinct tasks  [^1] The code and data are released at: https://github.com/YangLinyi/Supervised-Knowledge-Makes-Large-Language-Models-Better-In-context-Learners. Our empirical analysis sheds light on the advantages of incorporating discriminative models into LLMs and highlights the potential of our methodology in fostering more reliable LLMs.
@@ 3.1 Setup
Source models. As reported in GLUE-X (Yang et al., 2022), ELECTRA-large (Clark et al., 2020) achieves the best performance for both ID and OOD tasks over 21 small-scale pre-trained language models (maximum 774M parameters). Hence, we select ELECTRA-large as the SLM for NLU experiments, and RoBERTa-large (Liu et al., 2019) for QA. For evaluating the performance of SLM-enhanced LLMs, we select ChatGPT (OpenAI, 2023a) and Llama2-7B-chat (Touvron et al., 2023) as backbones, which are pre-trained on CommonCrawl, WebText, English Wiki, and others.
@@ 3.2 NLU Results
We also present the results of SuperContext with the prompt of the interpreter, which requires LLM to recall influential in-context examples and output rationales when making the predictions, indicating as SuperContext (+interpreter). To better understand the benefits of including the model confidence in the prompt, we present the results of SuperContext (w/o confidence). By comparing the results of SuperContext w/ and w/o confidence, we observe that including model confidence can bring significant improvements in the average performance for both ChatGPT and Llama2. Meanwhile, we find that for QNLI and QQP, Llama2 without the SLM’s confidence achieves the best performance among several methods. Our results also indicate that the interpreter can not bring significant benefits when compared to SuperContext in most of the tasks, except a slight improvement can be achieved on QNLI. It can be because the explain-then-predict prompt (Wang et al., 2022a) may not be suitable for incorporating with SuperContext, leading to information overload.
@@ 3.3 QA Results
Results of ChatGPT. The results are presented in Table 3. We find that although the original ChatGPT can achieve the highest accuracy for deterministic questions (79.35%), the exact match (EM) and accuracy for open questions can be significantly improved by SuperContext. In particular, the accuracy for no_answer questions can be improved from 32.35% (ChatGPT) to 54.65% (SuperContext), indicating the huge benefits. Besides, we find that even with the careful design of in-context learning prompts and filter methods, SuperContext still outperforms two in-context learning variants in terms of all metrics, indicating that pure in-context learning without fine-tuning LLMs brings no benefit to the QA task. Furthermore, SuperContext even outperforms the fine-tuned method in a multi-turn setting on all metrics. We believe that such a huge performance benefit (54.65% vs. 13.22%) compared to the traditional 16-shot in-context method when answering no_answer questions (“ACC.NO.”) proves that results achieved by discriminative models are effective enough to reduce the hallucination.
@@ 4.2 Interpretation Analysis
Learning from In-context Demonstrations. We explore how language models use long contexts. Figure 3 shows the influence of demonstrations during the inference stage, where the y-axis indicates how many times ChatGPT and InstructGPT take the $i_{th}$ in-context example as the emphasized one towards the prediction. The x-axis is sorted by the order of occurrence of in-context examples over 8 natural language understanding tasks. As shown in the figure, both ChatGPT and InstructGPT show a significant occurrence times difference among in-context examples. In particular, ChatGPT with 16-shot examples shows a trend of decreasing attention with the order of appearance. For example, the second in-context example has been paid attention to over 35,000 times while the last example only receives around 5,000 times attention. In terms of InstructGPT, we observe distinctive U-shaped occurrence times, which can be visualized in Figure 3(b). We find that the model tends to pay attention to the beginning or the end of the input context (in-context examples), and the attention significantly degrades in the middle of long contexts. This observation is consistent with the findings of (Liu et al., 2023) on the use of long contexts when performing downstream tasks, which suggests that model performance significantly degrades when models must access relevant information in the middle of long contexts and provide a new perspective for future long-context models. We also collect the sentence-level rationale generated by LLMs when making predictions, and count for the word frequency for each task of GLUE-X based on ChatGPT, aiming to provide the internal causes of OOD generalizability. However, the rationale is generated by LLMs and thus may contain hallucinations, which should be treated with caution and just for reference.
@@ 4.3 The Effect of SLM Confidence
Since we rely on the complementarity between SLMs and PLMs, SLMs must convey its certainly in task knowledge and uncertainly in domain knowledge to PLMs. The confidence score in the design serves a crucial role in such communication channels. We show the correlation between the confidence of SLMs and the prediction performance of LLMs. As shown in Figure 4, both ChatGPT and Llama2-7B-chat demonstrate a positive correlation between SLMs’ confidence and LLM’ performance, representing a high consistency between those models. The x-axis represents the confidence interval covering from 0.4-1.0, for example, 0.5 indicates the instances with the prediction confidence between 0.4-0.5. It is noteworthy that the confidence is computed by the zero-shot test based on SLMs trained on unseen domains, which indicates that high confidence requires a decent generalization ability of small models. We speculate that SuperContext shows superior performance than both SLMs and LLMs since it leverages the benefits of high consistency in discriminative models and the complementarity property of recent generative models. Besides, such a positive calibration law underscores the importance of involving both prediction and confidence in the prompt design of SuperContext. The data statistic of data quantity shows that most instances included in GLUE-X receive the highest confidence interval from 0.9 to 1.0, and this part of the data can be predicted with significantly higher accuracy than others. By comparing the experimental results of GPT-3.5 and Llama2-7B-chat, we find that when the confidence is more than 0.6, the average performance of GPT-3.5 is substantially better than Llama2-7B-chat.
@@ 5 Related Work
There is a line of work for improving the in-context learning performance by either constructing demonstrations (Arora et al., 2022; Si et al., 2022; Lyu et al., 2022; Gu et al., 2023; Ye et al., 2023; Dhuliawala et al., 2023) or framing an exploration of example selection methods (Wu et al., 2023; Wang et al., 2023b; Sun et al., 2023a; Agrawal et al., 2022; Wang et al., 2022c; b; Lu et al., 2022; Wang et al., 2023c) and even order (Lu et al., 2021; Zhao et al., 2021; Liu et al., 2021; 2023). The contrastive decoding method (Li et al., 2023b) considers the assistance smaller language model but requires external computation. Differently, SuperContext demonstrates its superior performance on OOD test data in a cost-effective manner.
@@
Knowledge in Context. Using external knowledge as auxiliary information to assist LLMs in providing truthful and timely responses represents an emerging solution (Mialon et al., 2023; Xiao et al., 2023) in recent. Traditional retrieve-based methods (Rubin et al., 2021; Ni et al., 2021; King & Flanigan, 2023) require a knowledge retriever as the prior step for guiding the generation of responses. Besides, the external knowledge source could extend beyond local documents to encompass the entire Internet (Ni et al., 2021; Gao et al., 2023). In addition, LLMs can leverage special plug-ins to improve their capabilities, such as Toolformer (Schick et al., 2023) and LangChain (Chase, 2022) for calling external APIs, and HuggingGPT (Shen et al., 2023) for using models.
Previous work either relies on web information and search engines for gaining external knowledge (Yu et al., 2023) or accomplishes planning tasks outside the NLP scope. (Xu et al., 2023a) evaluates the efficacy of small language models as plug-ins under an in-domain setting using GLUE and lacks an interpretation part to explain the reasons. SuperContext shares a conceptual similarity with SuperICL (Xu et al., 2023a) and HuggingGPT (Shen et al., 2023) in leveraging language model architectures. However, the key distinction lies in our approach’s application and analysis under out-of-distribution (OOD) conditions, a less explored area in the existing literature.
@@ 6 Conclusion and Future Work
Acknowledgement
We would like to thank the anonymous reviewers for their insightful comments and suggestions to help improve the paper. This publication has emanated from research conducted with the financial support of the Pioneer and “Leading Goose” R&D Program of Zhejiang under Grant Number 2022SDXHDX0003 and the National Natural Science Foundation of China Key Program under Grant Number 62336006.
@@ Appendix C Additional Results of Calibration Laws
We supplement the results of the calibration laws with two additional model groups, namely ELECTRA-large and InstrutGPT, in Figure 5. Consistent with ChatGPT and Llama2-7B-chat, both ELECTRA-large and InstructGPT exhibit a positive correlation between confidence and performance. Distinctively, the curve for InstrutGPT demonstrates more pronounced fluctuations, especially when the confidence is relatively low.